OpenStudy (anonymous):

solve the following pairs of simultaneous equation: x+y=7, x^2+y^2=25

OpenStudy (anonymous):

can somebody show me how to solve this? I did it, but when i checked it's wrong

OpenStudy (diyadiya):

x+y=7 therefore x = 7-y Replace x with 7-y in the equation \(x^2+y^2 = 25\) can you do that?

OpenStudy (diyadiya):

\[(x)^2+(y)^2=25\]\[(7-y)^2+(y)^2=25\]

OpenStudy (diyadiya):

@nnpatricia is this clear?

OpenStudy (anonymous):

yes! i've done that. making y=7-x. then later it leads to this 2x^2-14x-74. do you know how to factorize them?

OpenStudy (diyadiya):

I guess you did a small mistake there \[x^2 + (7-x)^2 =25\]\[x^2 + x^2 -14x + 49 =25\]\[2x^2 -14x + 24 =0\]

OpenStudy (diyadiya):

did you understand where you went wrong?

OpenStudy (phi):

(7-x)^2 or (7-x)(7-x) use FOIL First 7*7= 49 OUTER 7*-x= -7x INNER -x*7 = -7x LAST -x*-x= x^2 add up the four terms to get x^2-14x+49 use this in your original equation x^2 +(7-x)^2 = 25 (replace (7-x)^2 with the expanded form) x^2 + x^2-14x+49= 25 simplify by subtracting 25 from both sides of the equation add the 2 x^2 terms: 2x^2-14x+24=0 notice that all the terms are even, so simplify by dividing by 2: x^2-7x+12= 0/2 or just x^2-7x+12=0

OpenStudy (anonymous):

@Diyadiya yes! thanks :)

OpenStudy (phi):

the plus sign on the 3rd term +12 means both factors have the same sign the minus sign on the 2nd term -7x means the factors are negative

OpenStudy (anonymous):

@phi thanks for the very clear explanation too :)

OpenStudy (phi):

to factor, list the factors of the last term +12: 1,12 3,4 do any add up to 7? yes!